A constant current was passed through a solution of AuCl4
-
ions between gold electrodes. After a period of 10.00 minutes the cathode increased in weight by 1.314grams.
How much charge was passed and what was the current ?
1
Expert's answer
2019-08-05T02:08:49-0400

Solution.

We write the Faraday law for electrolysis:


"m = \frac{Q \times M}{F \times z}"

"Q = I \times \Delta t"

"Q = \frac{m \times F \times z}{M}"

F = 96484 C/mole

z = +3, because the charge of the gold atom in this ion is +3.

M(Au) = 196.967 g/mole

Q = 1930.98 C


"I = \frac{Q}{\Delta t}"

I = 3.22 A

Answer:

Q = 1930.98 C

I = 3.22 A